A circular pond 26 yd in diameter is surrounded by a gravel path 2 yd wide. the path is to be replaced by a brick walk costing $50 per square yard. How much will the walk cost?
Answer by rfer(16322)

A=3.14*13^2
A=530.66
-----------
A=3.14*15^2
A=706.5
---------
706.5-530.66=175.84 sq yrds path
175.84*50=$8,792